OpenStudy (konradzuse):

integral of x/sqrt(1-x^2) from (-1/8) to 0

OpenStudy (konradzuse):

\[\int\limits_{{-1/8}}^{0} \frac{x}{\sqrt{1-x^2}}\]

OpenStudy (konradzuse):

now normally it should be arcsin, but that x is on top, we cannot throw it infront of the integral though right?

OpenStudy (mimi_x3):

you can either use u-sub or trig-sub. which one do you want to use?

OpenStudy (konradzuse):

u sub i guess..

OpenStudy (konradzuse):

u = x du = 1

OpenStudy (konradzuse):

u on top? doesn't work...

OpenStudy (mimi_x3):

no..the bottom is \(1-x^2\) then let u =1-x^2 => du/dx = -2x

OpenStudy (konradzuse):

I figured it would be trig but yeah that works too. -1/2du = xdx

OpenStudy (konradzuse):

so then we have a u'/u

OpenStudy (mimi_x3):

you dont change it to -1/2du=xdx it wont work

OpenStudy (konradzuse):

why';s that? Don't we want to get the x on top as our u'?

OpenStudy (konradzuse):

hmm I guess that doesnt' work with the sqrt..

OpenStudy (mimi_x3):

\[\int\limits\frac{x}{\sqrt{1-x^{2}}} dx \] \[\int\limits\frac{x}{\sqrt{u}} *-\frac{du}{-2x} => \int\limits\frac{\cancel{x}}{\sqrt{u}} *-\frac{du}{\cancel{-2x}} \]

OpenStudy (mimi_x3):

do you understand? you have to cancel the \(x\)

OpenStudy (mimi_x3):

which leads to \[\frac{1}{2} \int\limits\frac{1}{\sqrt{u}} du\]

OpenStudy (konradzuse):

hmm never have seen that as du/something.

OpenStudy (mimi_x3):

then trig sub? lol

OpenStudy (mimi_x3):

i gotta go sorry

OpenStudy (konradzuse):

it's fine, thanks for the help thus far.

OpenStudy (konradzuse):

I guess when it's du = something dx then it's du/ that something?

OpenStudy (konradzuse):

or actually it's like I was showing before except instead of 1/2du it's 1/2x du..I got it.